# CosForce: A Force-Based General Model for Simulating Pedestrian Anticipation and Reaction Mechanisms

### v1

@@ 3.1 Model Rules
The self-driving force is driven by the free flow velocity vector $\boldsymbol{v}_{\text{free}}$, expressed as:
$\boldsymbol{f}_{i}={{m}_{i}}\left(\boldsymbol{v}_{\text{free}}-\boldsymbol{v}_{i}\right)/\tau$ (2)
Here, $\tau$ represents the relaxation time, and $m_{i}$ denotes the mass of pedestrian $i$, $m_{i}/\tau$ can be regarded as a fixed proportional coefficient.
The interaction force between a pedestrian and the nearest pedestrian is represented by the repulsive force, governed by a power-law decay function with respect to distance, expanded as follows:
${\boldsymbol{f}_{ij}}=\left({{c}_{1}}{{e}^{\left(\alpha{{r}_{ij}}-\left\|{\boldsymbol{d}_{ij}}\right\|\right)/{{\lambda}_{1}}}}+{{c}_{2}}\left\|{\boldsymbol{d}_{ij}}\right\|^{\eta}\left(1+cos\theta\right)\right){\boldsymbol{n}_{ij}}$ (3)
Here, $c_{1}$, $c_{2}$, $\lambda_{1}$, $\alpha$ (not a necessary coefficient), and $\eta$ are proportional coefficients.The angle $\theta$ is the angle between the relative distance and the relative velocity, $\theta=\angle\left(\boldsymbol{v}_{ij},\boldsymbol{d}_{ij}\right)\in\left[0,\pi\right]$. The symbols are defined as follows: $r_{ij}=r_{i}+r_{j}$, $\boldsymbol{v}_{ij}=\boldsymbol{v}_{i}-\boldsymbol{v}_{j}$, and $\boldsymbol{d}_{ij}=\boldsymbol{x}_{j}-\boldsymbol{x}_{i}$, ${\boldsymbol{n}_{ij}}=-{\boldsymbol{d}_{ij}}/\left\|{{\boldsymbol{d}_{ij}}}\right\|$.
The first term in the equation represents the contact force (elastic volume exclusion principle, exponentially decaying with distance), while the second term represents the non-contact force (repulsive effect, decaying with a power-law of distance). Given that, in low-density scenarios, the movement of pedestrian is also influenced by nearest neighbors from the far side (Murakami et al., 2021), a power-law function was employed rather than an exponential function for the calculation.
The interaction force between pedestrians and global obstacles is composed of two exponentially decaying terms with respect to distance, expanded as:
${\boldsymbol{f}_{iw}}=\left({{c}_{1}}{{e}^{\left(\alpha{{r}_{i}}-\left\|{\boldsymbol{d}_{iw}}\right\|\right)/{{\lambda}_{1}}}}+{{c}_{3}}{{e}^{\left(\alpha{{r}_{i}}-\left\|{\boldsymbol{d}_{iw}}\right\|\right)/{{\lambda}_{2}}}}\right){\boldsymbol{n}_{iw}}$ (4)
Here, ${{c}_{3}}$ and ${{\lambda}_{2}}\left({{\lambda}_{2}}\gg{{\lambda}_{1}}\right)$ represent proportionality coefficients, where $\left\|{\boldsymbol{d}_{iw}}\right\|$ denotes the perpendicular distance between pedestrian $i$ and the wall, and ${\boldsymbol{n}_{iw}}$ represents the standard vector opposite to the perpendicular line between pedestrian $i$ and the wall, ${\boldsymbol{n}_{iw}}=-{\boldsymbol{d}_{iw}}/\left\|{{\boldsymbol{d}_{iw}}}\right\|$.
The first term in the equation models the elastic volume exclusion principle exerted by the wall on the pedestrian under contact conditions, while the second term simulates the repulsive force of the wall on the pedestrian under non-contact conditions. If the virtual repulsive effect of the wall on the pedestrian is not required, the second term can be omitted. In this case, the wall can be considered a stationary obstacle composed of pedestrians.
3.2 Modeling Anticipation and Reaction Mechanisms via Cosine Functions
In this context, the anticipation mechanism of pedestrians is referred to as a cooperative game process among conflicting pedestrians (inducing phase separation). Conversely, the reaction mechanism is described as the convergence process among co-directional pedestrians (inducing aggregation).
The nearest-neighbor interaction force $\boldsymbol{f}_{ij}$ is adjusted nonlinearly using a cosine function of the angle $\theta$. When $\theta\to 0$, it indicates a potential collision trend, and $\boldsymbol{f}_{ij}$ increases. When $\theta\to\pi$, it indicates a separation trend between pedestrians, and $\boldsymbol{f}_{ij}$ decreases. With the introduction of the cosine function, the model approximately simulates the phase separation phenomenon between opposing pedestrians.
Phase separation of conflicting flows
Fig. 3(a) shows two agents, $i$ and $j$, in a cross-flow scenario with a potential collision, subject to mutually repulsive forces. Since agent $j$ is closer to the conflict point, the velocity reduction due to the repulsive force is smaller $\left(\left\|\boldsymbol{f}_{ji,y}\right\|<\left\|\boldsymbol{f}_{ij,x}\right\|\right)$. This trend is gradually amplified with the update of their relative positions, ultimately leading to agent $i$ yielding to agent $j$ (cooperative game). Strictly speaking, if the agents’ radii are ignored, non-cooperative games occur only when $\theta=0$.
Aggregation of co-directional flows
Fig. 3(b) illustrates that in unidirectional flow, agent $i$ is subjected to the repulsive force from agent $j$. If $\left\|\boldsymbol{v}_{j}\right\|>\left\|\boldsymbol{v}_{i}\right\|$, $\boldsymbol{f}_{ij}$ increases, and $\boldsymbol{v}_{i}$ decreases. If $\left\|\boldsymbol{v}_{j}\right\|<\left\|\boldsymbol{v}_{i}\right\|$, $\boldsymbol{f}_{ij}$ decreases, and $\boldsymbol{v}_{i}$ increases (considering the self-driving force). This leads to the aggregation of pedestrians in the same direction. Meanwhile, it satisfies the speed-density relationship under the constraints of the fundamental diagram.
4 Analytical Analysis
In the model, the force term is constrained by the exponential decay function and the power law decay function based on distance. In this section, a sensitivity analysis of the proportional coefficients governing the force term is performed.
4.1 Exponential Decay Term of Force
Fig. 4 illustrates the influence of the proportional coefficients $\lambda$ and $\alpha$ on the force term. A smaller $\lambda$ indicates that agents approach a rigid particle core, which may lead to a steep increase in interaction forces between pedestrians under high-density conditions, resulting in unrealistic instantaneous speeds. A larger value of $\lambda$ is indicative of pedestrian behavior resembling a collision with an elastic body; however, an excessively large $\lambda$ may result in an unrealistic overlap of body regions. The proportional coefficient $\alpha$ is used to adjust pedestrians’ dominance over nearby space. For a given decay coefficient ($\lambda$), increasing $\alpha$ enhances a pedestrian’s control over the neighboring area. Introducing $\alpha$ aims to simplify the setting of the coefficient $\lambda$. For an appropriate value of $\lambda$, the coefficient $\alpha$ can be removed from the model.
Figure 4: Sensitivity analysis of the nonlinear proportionality constants $\lambda$ (a) and $\alpha$ (b).
4.2 Power Law Decay Term of Force
Fig. 5 presents the influence of the proportional coefficient $\eta$ and angle $\theta$ on the power law decay force term. An increase in $\eta$ causes an increase in the decay rate of the force term. The parameter $\theta$ ensures that, for any given coefficient, when $\left\|{\boldsymbol{d}_{ij}}\right\|>{r}_{ij}$, ${f}_{ij}\in\left[0,2\right]$.
Figure 5: Sensitivity analysis of the nonlinear proportionality constant $\eta$ (a) and the angle $\theta$ (b).
The symbols setting are shown in Tab. 1.
Table 1: Symbol Descriptions and Reference Values
Symbol Description Reference value
@@ 4.2 Power Law Decay Term of Force
$v_{\rm{free}}$ Free-flow speed 1.2 m/s
@@
$\tau$ Relaxation time 1 s
$c_{1},c_{2},c_{3}$ Proportionality coefficients (linear) \
$\lambda_{1},\lambda_{2},\alpha,\eta$ Proportionality coefficients (nonlinear) \
@@ 5 Conclusions
In this paper, we proposed a force-based general model: CosForce. In the model, we employed forces constrained by cosine functions to model pedestrians’ Anticipation mechanism (separation between conflicting pedestrians) and reaction mechanism (aggregation of co-directional flow). Despite its simplicity, this model demonstrates unexpectedly performance, capable of simulating diverse pedestrian self-organizing phenomena, such as lane formation, stripe formation, and cross-channel formation.
Currently, the quantitative evaluation of the model is ongoing. Demonstration videos and codes are available in the Data Availability section. The videos present rough simulation results of the model.
@@
The codes and video of CosForce model can be found here: https://drive.google.com/drive/folders/1NYVnRp0z8VPuskfezMr51gB-sraOf6Iq?usp=drive_link (Google Drive).

### v2

@@ 3.3 Model Rules
The self-driving force is driven by the maximum velocity vector $\boldsymbol{v}_{\text{max}}$, expressed as:
$\boldsymbol{f}_{i}=\frac{{{m_{i}}}}{\tau}\left({\boldsymbol{v}_{\max}-{v}_{i}}\right)$ (11)
Here, $\tau$ represents the relaxation time, and $m_{i}$ denotes the mass of pedestrian $i$.
The interaction force between a pedestrian and the nearest pedestrian is represented by the repulsive force, which is expanded in both piecewise and continuous forms as Eqs. 12 and 13:
Piecewise function:
$\boldsymbol{f}_{ij}=\frac{{{m_{i}}}}{\tau}\left(\left\|{\boldsymbol{v}_{\max}}\right\|-\max\left({\min\left({\frac{{\left\|{\boldsymbol{d}_{ij}}\right\|-{r_{ij}}}}{{{t_{h}}}},\left\|{\boldsymbol{v}_{\max}}\right\|}\right),0}\right)\right){\left({1+\alpha\cos\theta}\right)}\boldsymbol{n}_{ij}$ (12)
Continuous function:
$\boldsymbol{f}_{ij}=\frac{{m_{i}}}{\tau}\left(\|\boldsymbol{v}_{\max}\|\left(1-\tanh\left(\frac{\|\boldsymbol{d}_{ij}\|-r_{ij}}{t_{h}}\right)\right)\right)\left(1+\alpha\cos\theta\right)\boldsymbol{n}_{ij}$ (13)
Here, The angle $\theta$ is the angle between the relative distance and the relative velocity, $\theta=\angle\left(\boldsymbol{v}_{ij},\boldsymbol{d}_{ij}\right)\in\left[0,\pi\right]$. Other symbols are defined as follows: $r_{ij}=r_{i}+r_{j}$, $\boldsymbol{v}_{ij}=\boldsymbol{v}_{i}-\boldsymbol{v}_{j}$, and $\boldsymbol{d}_{ij}=\boldsymbol{x}_{j}-\boldsymbol{x}_{i}$, ${\boldsymbol{n}_{ij}}=-{\boldsymbol{d}_{ij}}/\left\|{{\boldsymbol{d}_{ij}}}\right\|$. The simulation utilizes the difference calculation, with the nearest-neighbor repulsive forces constructed using a piecewise function.
In this process, obstacles are modeled as stationary pedestrians and are therefore included in the calculation of the repulsive force without requiring any additional rules. Different from the point-to-point interaction between pedestrians, the interaction between pedestrians and obstacles follows a point-to-line relationship. To account for all potential collision scenarios, the attentional eccentricity angle $\phi$ is set to a constant value of $\pi/2$.
The collision force between pedestrians is described by an exponential decay term, which is activated when the condition $\|\boldsymbol{d}_{ij}\|<r_{ij}$ is satisfied, as follows:
$\boldsymbol{f}_{c}=e^{\frac{r_{ij}-\left\|\boldsymbol{d}_{ij}\right\|}{\lambda}}\boldsymbol{n}_{ij}$ (14)
The parameter settings are presented in Tab. 1, where $m=20$ kg, $r=0.2$ m, $\tau=0.5$ s, $t_{h}=1.3$ s, and $\lambda=0.8$ are global parameters that remain constant across all simulations. The parameters $v_{\rm{max}}$, $\phi$, and $\alpha$ will be adjusted as required. The frequency for all simulations is set as 30 FPS, corresponding to a difference time of $1/30$ s.
Table 1: Parameter Descriptions and Reference Values
Parameter Description Reference value
@@
$v_{\rm{max}}$ Maximum speed 1.4 m/s
@@
$\tau$ Relaxation time 0.5 s
$t_{h}$ Approximate time headway 1.3 s
$\phi$ Attentional eccentricity angle [0 - $\pi$] rad
$\lambda$ Dimensionless coefficient 0.8/0.02 (scale in centimeters/meters)
$\alpha$ Dimensionless coefficient [0-1]
3.4 1-D dynamics of CosForce model
Approximated 2-D FVD Model under Force-based Modeling Framework
We have presented the two-dimensional form of the CosForce model to facilitate the analysis of the model’s properties. The one-dimensional equation of the CosForce model is expressed as:
$f=m_{i}a=\frac{m_{i}}{\tau}\left(\max\left(\min\left(\frac{{d}_{ij}-r_{ij}}{t_{h}},{v}_{\max}\right),0\right)-{v}\right)\left(1+\alpha\,\operatorname{sgn}\left({v}_{ij}\right)\right)$ (15)
It is clear that the one-dimensional CosForce model is very similar to the FVD model (Jiang et al., 2001). The difference is that the proportional coefficient does not undergo the same scale transformation as the velocity difference. For the sake of formal consistency, $1+\alpha\cos\theta$ can be expressed as $1+\alpha\frac{\|\boldsymbol{v}_{ij}\cos\theta\|}{\|\boldsymbol{v}_{\max}\|}$. Thus, the one-dimensional dynamical form is reformulated from $1+\alpha\,\operatorname{sgn}\left(v_{ij}\right)$ to the form of $1+\alpha\frac{v_{ij}}{v_{\max}}$. Such modifications, while increasing the complexity of the rules, were not found to lead to the anticipated improvement in model performance. The main issue lies in the inconsistency properties of speed differences between pedestrians and traffic: in traffic, a small speed difference is indicative of a stable flow state, whereas in crowds, a small speed difference may be interpreted as collective deceleration due to conflicts.
2-D OV model, when $\alpha$ = 0
When $\alpha=0$, the one-dimensional equation is given by Eq. 16, which is consistent with the OV model (Bando et al., 1995), the only difference being the empirical formula for the desired velocity.
$f=m_{i}a=\frac{m_{i}}{\tau}\left(\max\left(\min\left(\frac{{d}_{ij}-r_{ij}}{t_{h}},{v}_{\max}\right),0\right)-{v}\right)$ (16)
4 Numerical Validation
[…]
4.1 single-file pedestrians
First, we simulated the single-file motion in a channel under periodic boundary conditions, as shown in Fig. 6. Due to $\alpha=0$, the model behaves identically to the OV model. We set $N$=10, 20, 30, and 40 in the simulation to observe the dynamics of single-file pedestrians under different densities. The pre-simulation time was 900 steps, equivalent to 30 s, and simulation data for 60 s was obtained after the pre-simulation.
From the time-space diagram in Fig. 6, it can be observed that, as the global density increases, the pedestrian flow transitions through free flow, synchronized flow, and stop-and-go phases, and the traffic flow states remain highly stable.
Figure 6: Simulation of single-file pedestrians: (a) simulation snapshot, (b)-(d) time-space diagram at different densities.
4.2 unidirectional flow
Similarly, we conducted simulations of pedestrian unidirectional flow scenarios in a channel with periodic boundary conditions, as shown in Fig. 7. In the simulations, we incrementally set the number of pedestrians to 10 per simulation, conducting a total of 16 simulations covering scenarios from $N=10$ to $N=160$. The pre-simulation time was set to 900 time steps, corresponding to 30 s, and simulation data was obtained for 100 time steps (approximately 3.33 seconds) after the pre-simulation. Fig. 7 illustrates the state constraints of unidirectional pedestrian flow, including density versus speed, headway versus speed, and angular velocity versus speed. Due to the consistent sampling time, data obtained from high-density simulations are significantly more abundant than those from low-density simulations, leading to an uneven distribution of data in the plots.
Figure 7: Simulation of uni-directional pedestrian flow, (c)-(d) show the constraint relationships of the basic properties.
Density versus Speed
Fig. 7(c) presents the relationship between pedestrian speed and local density, which conforms to the characteristics of the pedestrian speed-density curve, showing an inverse proportional relationship in non-free-flow conditions.
Headway versus Speed
Fig. 7(d) illustrates the relationship between pedestrian speed and the nearest neighbor relative distance, with the distribution trend aligning with empirical results. Specifically, under non-free-flow conditions, speed exhibits a linear relationship with the nearest neighbor distance. The model is constructed based on a linear space-speed constraint, this result hence reasonably meets our expectations.
Angular velocity versus Speed
Fig. 7(e) shows the relationship between pedestrian angular velocity and speed, with the maximum angular velocity constraint exponent approximately equal to -1.3, slightly lower than our previous empirical observation (where the maximum angular velocity constraint exponent is -0.8, approximately).
5 "Phase separation" in crowds
In crowd dynamics, "phase separation" refers to the phenomenon where separation occurs between groups with different directions or states, induced by spontaneous order. Common examples include lane formation, stripe formation, and cross-channel formation. However, effective metrics for quantitatively estimating self-organization phenomena in crowds are still limited. Based on the mechanisms of self-organization: entropy reduction and steady state, we propose 3 quantitative analytical metrics: the average normalized velocity $\langle v\rangle$, the variance of normalized speed $Var(v)$, and the entropy of normalized speed $H$, with their formulas provided in Eqs. 17-19. These metrics provide quantitative standards for the efficiency of crowd movement, the volatility of crowd speed, and the degree of chaos in the crowd, respectively.
$\langle v\rangle=\sum\limits_{i\in N}\frac{\|\boldsymbol{v}_{i}\|}{N{v}_{\max}}$ (17)
$Var(v)=Var\left(\frac{\|\boldsymbol{v}_{i\in N}\|}{{v}_{\max}}\right)$ (18)
$H=-\sum_{j=1}^{10}p(\langle{v}_{i\in N}\rangle_{j})\log(p(\langle{v}_{i\in N}\rangle_{j}))$ (19)
The simulation environment is set as an 8m $\times$ 8m with fully periodic boundary conditions. Snapshots of the three scenarios are shown in Fig. 8. In the initial state, all pedestrians are uniformly distributed, with an initial velocity of 0. The simulation time for a single scenario is 3000 steps (100 seconds).
Figure 8: Simulation snapshots of pedestrian "phase separation": (a) lane formation, (b) stripe formation, (c) cross-channel formation.
5.1 Lane formation
[…]
Figure 9: Variation of crowd velocity metric over time: (a) average normalized speed, (b) variance of normalized speed, (d) entropy of normalized speed.
Fig. 9 shows that within a short initial period (approximately 30 seconds), the average normalized speed increases and reaches a stable state, maintaining an average value of approximately 0.6 m/s. Correspondingly, the variance and entropy of the normalized speeds are also observed to decrease and stabilize.
5.2 Stripe formation
[…]
Fig. 10 presents a trend similar to that in Fig. 9, i.e., the decrease in the average speed is accompanied by a reduction in both the speed variance and entropy. In terms of the changes in the mean, this trend is observed to be quite steady. Such properties indicate the process of spontaneous order formation within the system, and this formation of order typically occurs within a time frame ranging from several seconds to a few tens of seconds.
Figure 10: Variation of crowd velocity metric over time: (a) average normalized speed, (b) variance of normalized speed, (d) entropy of normalized speed.
5.3 Cross-channel formation
[…]
Figure 11: Variation of crowd velocity metric over time: (a) average normalized speed, (b) variance of normalized speed, (d) entropy of normalized speed.
10 independent simulations were conducted corresponding to different dynamic pedestrian ratios. The resulting trends of the average normalized speed, as well as the variance and entropy of normalized speed, are shown in Fig. 11. The configuration with a 100% dynamic pedestrian ratio can be considered as the control group, representing the unified unidirectional flow of the crowd.
From Fig. 11, we can still observe similar self-organization trends, namely, the crowd’s speed steadily increases, while the speed variance and speed entropy show a stable decreasing trend.
6 "Catfish effect" in crowds
The continuous dynamics of crowds are one of the fundamental causes of crowd crush. Some basic characteristics of crowd stability have been empirically observed, with contributions from pioneering work (Gu et al., 2024; Ma et al., 2013; Echeverría-Huarte et al., 2022) on the oscillation phenomenon within crowds. The sensitivity of pedestrians to spatial variations was revealed in our previous work. In general, people seem reluctant to stop, indicating that pedestrians strive to maintain dynamic, resulting in low velocities and high angular velocities.
Figure 12: Simulation of the "catfish effect" in crowds: (a) physical illustration of the background crowd, (b) simulation snapshots when the number of active pedestrians is 0, 1, and 10, (c) velocity field distribution when the number of active pedestrians is 0, 1, and 10, (d)-(f): variation of speed metric over time.
[…]
Overall, the background crowd consists of "passive pedestrians", providing the system’s background environment. Our aim is to study the changes in the system state caused by the "active pedestrians" and explore the potential phenomenon of the "catfish effect", which originates from our earlier conjecture (Wang et al., 2023). The simulation environment is set as an 8m $\times$ 8m with fully periodic boundary conditions. Snapshots of the three scenarios are shown in Fig. 8. In the initial state, all pedestrians are uniformly distributed, with an initial speed of 0 m/s. The simulation time for a single scenario is 3000 steps (100 s).
The background crowd is shown in Fig. 12(b) and exhibits the following characteristics:
A closed system (full periodic boundary conditions),
No friction,
Collisional forces are symmetric,
Rules lead to a structured distribution of pedestrians (asymmetric forces, being statistically symmetric, result in negligible external net forces).
These properties indicate that the system is close to momentum conservation. The trends of the orthogonal direction and the mean scalar velocity are shown in Fig. 13(a). Based on these properties, Fig. 12(a) provides an equivalent physical system for the background crowd, which is a closed system composed of four fixed points constrained by Hookean springs and balls, where the balls undergo uniform circular motion.
In the comparative scenarios, 1 and 10 "active pedestrians" were added, respectively. Fig. 12(c) shows the distribution of the velocity field, with the corresponding quantitative data presented in Figs. 12(d)-12(f). From the statistical data, after the introduction of polarized "active pedestrians", the system’s kinetic energy showed an increment, accompanied by a larger variance of speed. However, an intriguing phenomenon is that the entropy of velocity decreases. The underlying reasons remain unexplained. These results suggest that the presence of "active pedestrians" enhances the ordering of the dynamic crowd.
Figure 13: Average orthogonal velocity and scalar speed over time for the number of "active pedestrians": (a) $N$ = 0, (b) $N$=10.
From the perspective of momentum analysis, we examined the orthogonal velocity distribution when the number of "active pedestrians" was 0 and 10, respectively. In the control group, the velocity component in the orthogonal direction fluctuated around zero, indicating that the system was near equilibrium. In Fig. 13(b), the inclusion of "active pedestrians" led to a polarization effect in the $x$-direction, resulting in an increase in the average normalized speed in the $x$-direction by approximately 0.1. From the perspective of momentum variation, the momentum increment of the 10 "active pedestrians" did not exceed 10 units (assuming uniform pedestrian mass and a normalized maximum speed of 1), while the average velocity in the $x$-direction increased by approximately 0.1, corresponding to a total momentum gain of about 25 units in the system. The fluctuation trend in the $y$-direction velocity remained consistent with that of the control group.
Referring to Fig. 12(c), we observed that the momentum in different directions within the control group canceled each other out, leading to a small total momentum that fluctuated around zero. When the "active pedestrians" entered the crowd, adjacent individuals gradually adjusted their velocity, and the random momentum became more organized (i.e., polarized), resulting in a significant increase in the system’s total momentum (gain > increment).
7 Limitations
In the simulation, we also explored the limitations of the model, which mainly include two aspects:
Inability to simulate tactical-level behaviors: As shown in Fig. 13 (a) and (b), even simulating some local tactical-level behaviors, such as short-term route decisions, is beyond the capability of the model, which is a common limitation of most operational-level models.
Distortion in response to static environments: In certain static environments, pedestrians may become trapped even when space is available. Within the framework of force-based models, there exist equilibrium positions in static environments that cause the pedestrian’s force balance, resembling a black hole, where pedestrians passing nearby will drawn towards these equilibrium points.
Figure 14: Limitations of the model: comparison between experimental and simulation trajectories. (a) antipode experiment, (b) crowd-cross experiment (static).
@@ 8 Conclusions
In this paper, we propose a force-based general model, CosForce. The model consists of 8 parameters (6 agent state parameters and 2 dimensionless coefficients), along with four equations.
In order to simulate pedestrians’ anticipation and reaction behaviors, as well as the associated collision avoidance mechanisms, symmetric and asymmetric forces constrained by cosine functions are employed. Despite its simplicity, the model demonstrates effective performance, being capable of simulating a range of pedestrian self-organization phenomena, among others.
We explored "phase separation" and the "catfish effect" within crowds based on simulations of the CosForce model. First, several quantitative metrics were introduced to evaluate pedestrians’ self-organization process, including the average normalized speed, the variance of normalized speed, and the entropy of normalized speed. Given that the theory of pedestrian self-organization is well-established and supported by empirical observations, the simulation results align with our expectations. On the other hand, the simulation of the "catfish effect" within crowds is entirely theoretical. A key finding is that while "active pedestrians" contribute to an additional gain in system momentum, the system’s entropy of speed decreases, simultaneously. In our simulation, "active pedestrians" induce the formation of order within the crowd. The rules governing pedestrian motion are simple and reasonable, with the desired speed derived from an empirical equation that ensures dimensional consistency. We consider the results valid. In summary, further discoveries rely on additional exploration, particularly through empirical observations.
Theoretically, compared to traditional force-based models, the complexity of the CosForce model can be reduced from $O(n^{2})$ to $O(n)$. Currently, we are pursuing this via a structured calculation method.
@@
The code can be found at: https://drive.google.com/drive/folders/1NYVnRp0z8VPuskfezMr51gB-sraOf6Iq?usp=drive_link (Google Drive), and a short video summary is available at: https://www.bilibili.com/video/BV17B1oYVEQm/.